OpenStudy (love_to_love_you):

Which function is not a power function? f(x) = x f(x) = x^1/3 f(x) = 1/x f(x) = |x|

OpenStudy (jdoe0001):

any ideas?

OpenStudy (jdoe0001):

hint: \(\bf x^{\frac{1}{3}}\impliedby obvious \\ \quad \\ \cfrac{1}{x}\iff x^{-1}\impliedby \textit{also a power one} \\ \quad \\ |x|\iff +x,-x\iff \sqrt{x^2}\implies (x^2)^{\frac{1}{2}}\)

OpenStudy (love_to_love_you):

so |x|?

OpenStudy (jdoe0001):

hmmm well... nope obvously is a power function, since it uses power to get the +/- signs \(\bf x \iff x^1\) to be fair, however, a 1 itself, doesn't make it a power function so, it'd be "x"

OpenStudy (zarkon):

https://en.wikipedia.org/wiki/Power_function

OpenStudy (zarkon):

you are correct... \(|x|\) is not a power function